Use the "ac" test to determine if the following trinomial can be factored.
x^2+2x+3
Please show me the steps to determine this. Thank you!

Answer by stanbon(75887)

Use the "ac" test to determine if the following trinomial can be factored.
x^2+2x+3
--------------
Your Problem:
a = 1 ; b = 2 ; c = 3
Think of two numbers whose product is a*c = 1*3 = 3
and whose sum is b = 2
--------------------------
There are no such numbers.
The ac method cannot be used to factor your problem.
===========================
Cheers,
Stan H.